# Emerging Technologies for 6G Non-Terrestrial-Networks: From Academia to Industrial Applications

### v1

@@ VI-B3 Generative AI for NTN
Exploring the integration of Generative Artificial Intelligence (GAI) in NTN is a promising future research direction with huge potential. Generative AI, known for its ability to create synthetic data and generate novel content, can be utilized to address several challenges in NTN. For instance, it can be employed to generate high-quality synthetic data, complementing the training and testing of traditional AI techniques in the dynamic and diverse NTN environments. Moreover, GAI can contribute to enhance the security of NTN by generating realistic attack scenarios, enabling robust evaluation of network defense mechanisms. Another approach is utilizing GAI to generate network data for simulating various network conditions, thereby aiding network optimization and predicting potential bottlenecks.
VII CONCLUSION
@@
In this paper, we have presented a comprehensive survey of NTN for 6G networks from both academic and industry perspective. Particularly, we have presented the in-depth tutorial on NTN and the enabling technologies including network slicing, AI/ML, and ORAN. Then, we have surveyed state-of-the-art network slicing and AI/ML approaches that are proposed to address various challenges of NTN in the literature. Moreover, we have presented how ORAN has been utilized for NTN from the industry standpoint. Finally, we have discussed the current challenges as well as open issues and introduced promising technologies as future research directions of NTNs.

### v2

@@ 6.2.3 Generative AI for NTN
Exploring the integration of Generative Artificial Intelligence (GAI) in NTN is a promising future research direction with huge potential. Generative AI, known for its ability to create synthetic data and generate novel content [148], can be utilized to address several challenges in NTN. For instance, it can be employed to generate high-quality synthetic data, complementing the training and testing of traditional AI techniques in the dynamic and diverse NTN environments. Moreover, GAI can enhance the security of NTN by generating realistic attack scenarios, enabling robust evaluation of network defense mechanisms. Another approach is utilizing GAI to generate network data for simulating various network conditions, thereby aiding network optimization and predicting potential bottlenecks.
6.2.4 Holographic MIMO and Intelligent Reflecting Surfaces
The recent advancement of antenna technologies has opened promising research directions that can significantly enhance the performance of NTN. For example, holographic MIMO is an advanced antenna technology that leverages principles from holography to enable highly directional and focused transmissions from a single aperture. This technique enables highly directional transmissions while mitigating interference, thereby improving spectrum utilization and throughput for NTN systems [149]. Additionally, the concept of intelligent reflecting surfaces (IRS) has emerged as a promising enabling technology for NTNs. IRSs comprise arrays of reconfigurable meta-surfaces that can alter the propagation properties of impinging electromagnetic waves [150]. By strategically deploying IRSs in conjunction with satellites or aerial platforms, NTN providers can optimize signal reflections, extend coverage areas, and enhance the overall quality of service. Moreover, another promising direction could be the combination of IRS and holographic MIMO, e.g., deploying IRSs in strategic locations to intelligently reflect and focus the multiple beams generated by a holographic MIMO transmitter towards intended receivers.
7 CONCLUSION
@@
In this paper, we have presented a comprehensive survey of NTN for 6G networks from both academic and industry perspectives. Particularly, we have presented an in-depth tutorial on NTN and the enabling technologies including network slicing, AI/ML, and ORAN. Then, we have surveyed state-of-the-art network slicing and AI/ML approaches that are proposed to address various challenges of NTN in the literature. Moreover, we have presented how ORAN has been utilized for NTN from the industry standpoint. Finally, we have discussed the current challenges as well as open issues and introduced promising technologies as future research directions of NTNs.